OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8)

OpenStudy (anonymous):

we will find the slope by using the slope formula: slope(m) = (y2 - y1) / (x2 - x1) (4,5) x1 = 4 and y1 = 5 (6,8) x2 = 6 and y2 = 8 now we sub slope(m) = (8 - 5) / (6 - 4) m = 3/2 ===> your slope now we use y = mx + b form and solve for b (4,5) x = 4 and y = 5 m = 3/2 now we sub y = mx + b 5 = 3/2(4) + b 5 = 12/2 + b 5 = 6 + b 5 - 6 = b -1 = b so your line is : y = 3/2x - 1 ==> slope intercept form

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8).